From two points each on the opposite sides of a river . The angles of elevation of the top of an 80 feet tree are 60° and 30°. The points and the tree are in the same straight line , which is perpendicular to the river. How wide is the river ?

1
Expert's answer
2022-02-13T17:45:16-0500

"\tan\alpha=\frac h{x+a}," "\tan\beta=\frac ha,"

"x=\frac h{\tan \alpha}-a," "a=\frac h{\tan \beta},"

"x=h(\frac 1{\tan \alpha}-\frac 1{\tan\beta})=33.8~ft."